# Simple Policy Optimization

### v8

@@ 1 Introduction
Despite its success, the optimization behavior of PPO remains insufficiently understood. Although PPO aims to constrain the probability ratio deviations between successive policies, it often fails to keep these ratios within bounds (Ilyas et al., 2018; Engstrom et al., 2020; Wang et al., 2020). In some tasks, the ratios can even escalate to values as high as $40$ (Wang et al., 2020). Furthermore, studies have revealed that PPO’s performance is highly dependent on “code-level optimizations” (Andrychowicz et al., 2021; Huang et al., 2022a). The implementation of PPO includes numerous code-level details that critically influence its effectiveness (Engstrom et al., 2020; Huang et al., 2022b).
@@ 4 Methodology
where $\delta_{\mathrm{KL}}>0$ is a predefined threshold. Let $\delta_{\mathrm{TV}}=\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}$, we establish that $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$.
@@
For any given $\delta_{\mathrm{KL}}$ and $\tilde{\pi}\in\Omega_{\mathrm{KL}}$, using Pinsker’s inequality, we have $D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\leq\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\leq\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}=\delta_{\mathrm{TV}}$, therefore $\tilde{\pi}\in\Omega_{\mathrm{KL}}\Longrightarrow\tilde{\pi}\in\Omega_{\mathrm{TV}}$, which means $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$, concluding the proof. ∎
@@
be the lower bounds of performance improvement with TV divergence and KL divergence. Let $\delta_{\mathrm{TV}}=\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}$, denote
@@ 6.2 Scaling Policy Network
Figure 6: Training curves of SPO using ResNet-18 as the encoder compared to the PPO and SPO using original CNN. The mean and standard deviation are shown across 3 random seeds, and the red dashed line shows $\epsilon=0.2$.
@@
We also trained the ResNet-18111Since Bhatt et al. (2019) demonstrated that batch normalization is harmful to RL, we removed batch normalization and adjusted ResNet-18 for input and output from our implementation. (He et al., 2016) as the encoder on the Atari 2600 benchmark, the results are shown in Figure 6. As the network’s capacity increases, the performance of SPO is significantly improved. Moreover, SPO can still maintain a good probability ratio constraint, thereby benefiting from the theoretical lower bound (10). In contrast, it is challenging to train large neural networks with PPO because the probability ratio cannot be controlled during training, even employing a smaller $\epsilon=0.1$.
@@ Acknowledgements
We thank the anonymous ICLR and ICML reviewers for their insightful and constructive comments, as well as our collaborators for their valuable suggestions.
@@ Appendix A Atari 2600
Figure 8: The performance of SPO compared to PPO across 35 games in Atari 2600 environment.
Appendix B Hyperparameter Settings
@@ Appendix C More Results
Figure 10: Training performance of PPO and SPO in Atari 2600. The mean and standard deviation are shown across 3 random seeds.

### v9

@@ 1 Introduction
Despite its success, the optimization behavior of PPO remains insufficiently understood. Although PPO aims to constrain the probability ratio deviations between successive policies, it often fails to keep these ratios within bounds (Ilyas et al., 2018; Engstrom et al., 2020; Wang et al., 2020). In some tasks, the ratios can even escalate to values as high as 404040 (Wang et al., 2020). Furthermore, studies have revealed that PPO’s performance is highly dependent on “code-level optimizations” (Andrychowicz et al., 2021; Huang et al., 2022a). The implementation of PPO includes numerous code-level details that critically influence its effectiveness (Engstrom et al., 2020; Huang et al., 2022b).
@@ 4 Methodology
where $\delta_{\mathrm{KL}}>0$ is a predefined threshold. Let $\delta_{\mathrm{TV}}\geq\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}$, we establish that $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$.
@@
For any given $\delta_{\mathrm{KL}}$ and $\tilde{\pi}\in\Omega_{\mathrm{KL}}$, using Pinsker’s inequality, we have $D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\leq\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\leq\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}\leq\delta_{\mathrm{TV}}$, therefore $\tilde{\pi}\in\Omega_{\mathrm{KL}}\Longrightarrow\tilde{\pi}\in\Omega_{\mathrm{TV}}$, which means $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$, concluding the proof. ∎
@@
be the lower bounds of performance improvement with TV divergence and KL divergence. Let $\delta_{\mathrm{TV}}\geq\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}$, denote
@@ 6.2 Scaling Policy Network
Figure 6: Training performance of SPO using ResNet-18 as the encoder compared to the PPO and SPO using default CNN (shown with the reference line). The mean and standard deviation are shown across 3 random seeds, and the red dashed line represents $\epsilon=0.2$.
@@
We also trained the ResNet-18111Since Bhatt et al. (2019) demonstrated that batch normalization is harmful to RL training, we removed batch normalization. (He et al., 2016) as the encoder on the Atari 2600 benchmark, the results are shown in Figure 6. As the network’s capacity increases, the performance of SPO is significantly improved. Moreover, SPO can still maintain a good probability ratio constraint, thereby benefiting from the theoretical lower bound (10). In contrast, it is challenging to train large neural networks with PPO because the probability ratio cannot be controlled during training, even employing a smaller $\epsilon=0.1$.
@@ Acknowledgements
We would like to thank the anonymous ICLR and ICML reviewers for their insightful and constructive comments.
@@ Appendix A Atari 2600
Figure 8: Final performance of SPO compared to PPO across 35 games in Atari 2600 environment, using default CNN as the encoder.
Appendix B Hyperparameters
@@ Appendix C More Results
Figure 10: Training performance on Atari 2600 using ResNet-18 as the encoder, with a fixed learning rate of 0.0001. The mean and standard deviation are shown across 3 random seeds.